OpenStudy (anonymous):

What's the idea to solve for m the equation m(3^(2m)+3) = 0 mod 7 ?

OpenStudy (ikram002p):

3antarrr kefak

OpenStudy (ikram002p):

so2alak heek ? \(\large m(3^{2m}+3) = 0 \mod 7 \)

OpenStudy (anonymous):

Ahlan ikram, Exactly

OpenStudy (ikram002p):

what i can see is since 7 is prime humm let me think :O

OpenStudy (ikram002p):

ok 3antar 7 is prime 3ashan heek e7na 3ena two cases :- m=0 mod 7 or 3^(2m)+3 = 0 mod 7

OpenStudy (ikram002p):

case 1 m=7k

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Why the two cases @ikram002p ?

OpenStudy (ikram002p):

case 2 3^(2m)+3 = 0 mod 7 we need to show which m makes i true since 3 is prime then 3^2m mod 7=3 mod 7 so 3 mod 7 + 3 = 6 mod 7 \(\neq\) 0 mod 7 so the only case is m=2k

OpenStudy (ikram002p):

7 is prime , u cant write it as a multiple of two numbers

OpenStudy (anonymous):

Ah' ok

OpenStudy (anonymous):

@ikram002p What happened when 3 was prime ?

OpenStudy (anonymous):

@rational have you an idea ?

OpenStudy (rational):

case1 is trivial : m = 7k

OpenStudy (rational):

case2 : You need to solve \(3^{2m}+3 \equiv 0 \mod 7\)

OpenStudy (rational):

you familiar with Fermat little thm ?

OpenStudy (rational):

\(3^{2m}+3 \equiv 0 \mod 7\) \(\implies (3^2)^{m} \equiv 4 \mod 7\) \(\implies 2^{m} \equiv 4 \mod 7\) \(\implies 2^{m-2} \equiv 1 \mod 7\)

OpenStudy (rational):

see if that looks okay so far

OpenStudy (anonymous):

@rational Awesome go ahead

OpenStudy (anonymous):

are you going to use log base 2, :) ?

OpenStudy (rational):

haha much simpler than that

OpenStudy (rational):

one sec

OpenStudy (rational):

\(\implies 2^{m-2} \equiv 1 \mod 7\) By FLT, \(m-2 \equiv 0 \mod 6 \implies m = 6k+2\) gives u solutions, but these are not exhaustive. there will be other solutions as well.

OpenStudy (rational):

\(m = 7k, ~ 6k+2\) are part of solutions. let me think a bit to conclude on other solutions

OpenStudy (rational):

wil get back in the morning, this wil take time :)

OpenStudy (anonymous):

@rational How did you use the FLT ?

OpenStudy (rational):

\(\large a^{p-1} = 1 \mod p\)

OpenStudy (rational):

IF \(m-2 = 6\), then it will be a solution of : \(2^{m-2} = 1 \mod 7\), right ?

OpenStudy (anonymous):

What's p in our case ? 7, right ?

OpenStudy (anonymous):

= or equiv ?

OpenStudy (rational):

equiv

OpenStudy (rational):

and yes p=7